OpenStudy (anonymous):

HELP ASAP A sequence is defined by the recursive function f(n + 1) = f(n) – 2. If f(1) = 10, what is f(3)? 1 6 8 30

OpenStudy (welshfella):

f(2) = f(1) - 2 = 10 - 2 = 8 f(3) = ?